Reports 1-1 of 1 Clear search Modify search
AdV-INJ (Suspended injection bench (SIB1) body)
gosselin, melo, bersanetti, pinto - 23:41 Tuesday 12 September 2023 (61616) Print this report
SIB1 noise injections

We made some noise injections on SIB1 degrees of freedom in order to see the effect on the ITF.

Below is the GPS associated with each injection and the level of noise applied through python:

PR_X (ASC_noise_color_ON) to inject noise on SIB1_TY,
To be noted that we were not injecting noise on PR but on B2, the name PR_X has not been uptaded after the new alignment scheme.

  • 15:33:22 (1e-10)
  • 15:35:21 (1e-8)
  • 15:38:03 (1e-5)
  • 15:40:50 (5e-5)
  • 15:42:50 (1e-6)
  • 15:44:00 (1e-4)
  • 15:46:00 (5e-4)
  • 15:50:50 (1e-3)
  • 15:58:00 (5e-3) Here we could see some structures at 28.2 Hz in Hrec and some lines higher at 157 Hz, 225 Hz, 300 Hz, with coherence between ASC_noise vs. Hrec at 28.2 Hz and 78 Hz
  • 16:03:40 (1e-2) Coherence between 20 and 28 Hz
  • 16:07:40 (2e-2) Coherence between 20 and 40 Hz
  • 16:11:40 (4e-2) Coherence between 15 and 40 Hz
  • 16:17:20 Noise OFF

Just after having switched off the noise the ITF unlocked.

We could mainly see some structure appearing at low frequency, between 10 and 30 Hz (see figure 1).
 

PR_Y (ASC_noise_color_ON) to inject noise on SIB1_TX to inject noise on SIB1_TX.
To be noted that we were not injecting noise on PR but on B2, the name PR_Y has not been uptaded after the new alignment scheme.
 

  • 17:24:30 (1e-8)
  • 17:26:30 (1e-6)
  • 17:29:10 (1e-5)
  • 17:31:10 (1e-4)
  • 17:33:20 (1e-3)
  • 17:35:20 (5e-3)
  • 17:37:22 (1e-2)
  • 17:39:10 (2e-2) Coherence between 10 and 20 Hz
  • 17:45:30 (4e-2) peak seen at 8.5 Hz on SIB2_B2_QD1_8MHz_V_I/Q; coherence between 10 and 20 Hz
  • 17:52:33 Noise OFF
  • 17:55:20 (8e-2) Structure on DARM between 5 and 15 Hz, lines on Hrec at 228 Hz, 1150Hz, 1111 Hz, 2488 Hz

The ITF unlocked. Most likely due to the noise and the structure that we could see at low frequency. More injection at the end of the shift with less noise at low frequency.

IB_TX_SENSING_NOISE (Injections on ISYSnoise_External, changing the weight in the driving IB matrix)

  • 19:11:06 (0.01) 61 Hz line?
  • 19:15:10 (0.05)
  • 19:22:36 (0.1)
  • 19:27:51 (1)
  • 19:35:01 (2)
  • 19:38:08 (4)
  • 19:42:08 (10)
  • 19:45:53 (OFF)
  • 19:51:56 (10)
  • 19:55:57 (20)
  • 19:58:56 (100)
  • 20:04:04 (500) Nothing seen on DARM or Hrec
  • 20:09:15 (0 V)

No visible effect on Hrec (figure 3)

PR_Y with new filter to try to reduce the noise injected at lw frequency (see figure 4)

  • 20:14:10 (1e-6)
  • 20:19:05 (1e-2)
  • 20:22:35 (1e-6)
  • 20:43:40 (5e-6)

The ITF unlocked.

No big effects on Hrec (figure 2).

Images attached to this report
Comments to this report:
gosselin, derossi - 22:48 Tuesday 19 September 2023 (61714) Print this report

We redid the noise injection in LN3

DOF GPS start UTC duration filter amplitude
quiet 21:45:00 5 min    
PR X 22:08:20     5e-9
" 22:11:08 3 min   5e-8
quiet 22:15:00 2 min    
PR Y 22:22:20 2 min   5e-8
" 22:22:35 2 min   1e-7
  22:24:05     1e-6 too high, we unlocked

For PR_X (ie SIB1 TY), we were sending more noise at high frequency and less at low frequency than last week (see the filter at the bottom of the entry), we saw many new structures in the range 50-200 Hz (see figure 1 and 2).
For PR_Y we were sending noise similar to what we did last week but it made the interferometer unlock.

# ASC Noise filter butterworth
ACL_FILTER_SET             ASC_noise_butter        1    1    10  12
ACL_FILTER_BUTTERWORTH     ASC_noise_butter    "lowpass"  2 3    200
ACL_FILTER_BUTTERWORTH     ASC_noise_butter    "highpass" 10 3    50

Images attached to this comment
derossi - 14:10 Wednesday 20 September 2023 (61726) Print this report

two mistakes in the GPS of the injections of Sept. the 19th :

1) it is LT and not UTC

2) for PR TY with amplitude 5e-8 is starting at 22:20:20, for 2 minutes

derossi, melo, gosselin - 23:59 Monday 09 October 2023 (61947) Print this report

today we wanted to perform again injections on SIB1 with a new filter (in ASC_pre)

# ASC Noise filter butterworth
ACL_FILTER_SET             ASC_noise_butter        1    1    10  24
ACL_FILTER_BUTTERWORTH     ASC_noise_butter    "highpass" 10 3    100

However, we only made some trials with this filter on PR Y between 21:24:02 UTC and 21:25:21 UTC and we unlocked ( 1e-7 amplitude was too high ).

Prior to try this filter we inejcted white noise on PR X (by mistake, because we did not load the new filter). We report however the GPS:

- quiet 20:59:30 UTC 4 min

- 21:10:56 4 min ampl 1e-2 on PR X

- 21:16:00 2 min ampl 5e-2 on PR X (fig. 1)

Images attached to this comment
melo, ruggi - 10:18 Wednesday 13 December 2023 (62628) Print this report

With the data of Sept 12th, I have computed the noise projections for the TY and TZ DoF of the SIB1 bench.

In Fig. 1 we can see the coherence between the DoFs and Hrec. As can be seen, the coherence is higher between 15 and 40 Hz for TY and between 7 and 20 Hz for TZ.

The noise projection on Hrec, done considering the coherence window, are shown in Fig. 2. The noise is below the O4/10 sensitivity curve and we can notice a structure around 28.3 Hz that is excited when the noise is injected.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×